OpenStudy (anonymous):

What is 5 1/3 · (–4 4/5)?

OpenStudy (anonymous):

I am pretty sure that its -20 4/15.....

OpenStudy (anonymous):

turn them to improper fractions.. 5 1/3 = 16/3 -4 4/5 = - 24/5 now we multiply keeping in mind that a positive times a negative is a negative 16/3 * - 24/5 = -384/15 = - 25 3/5

OpenStudy (anonymous):

ok thank you so much :-)

OpenStudy (anonymous):

sure thing :)